Some acetone is poured into a thick-walled, 150 cm 3 test tube at 293 K and 100 kPa. The tube was then sealed and heated to 623 K. What is the maximum mass of acetone that can be used for this process if the pressure cannot exceed 5.00 MPa? The amount of air in the test tube can be neglected.


1
Expert's answer
2021-01-11T03:54:50-0500

p*V = n*R*T;

5MPa = 49.3462 atm

n = p*V/R/T = (49.3462 atm)*(0.15 L)/(0.082 L*atm/mol/K)/(623 K) = 0.1449 mol

m = n*Mr = (0.1449 mol)*(58.08 g/mol) = 8.4153 g